OpenStudy (anonymous):

find the probability of each event. a) rolling 2 dice and getting a sum of 7 or 11 b) tossing a coin 5 times and getting a tail everytime. c) choosing a queen or a red ace from a deck of cards 3:36 pm

OpenStudy (anonymous):

2/9, 1/32, 5/56

OpenStudy (anonymous):

a) 5/36 + 2/36 = 7/36 b) (1/2)^5 = 1/32 c) 4/52 + 2/52 = 6/52 = 3/26 (assuming no jokers)

OpenStudy (anonymous):

hmm i'm a bit rubbish at explanations, but i'll give it a go .. theoretical probability is a probability / a chance of something that you can't test, the opposite from the other one, for example, the probability of goin thro a black hole and entering another universe would be a theoretical prob as you can't test it .. something like tht i believe